Apply your knowledge on ratio and share R25 between Thando and Sibi in the ratio 2:3.


1
Expert's answer
2021-08-10T09:17:57-0400

Thando's share: "25*\frac{2}{2+3}=R10".

Sibi's share: "25*\frac{3}{2+3}=R15."